# Identity-GRPO: Optimizing Multi-Human Identity-preserving Video Generation via Reinforcement Learning

### v2

@@ 4.2 GRPO Training
Experiment Setup. We utilized the reference images from the preference dataset we constructed, along with the corresponding prompts generated by our prompt finetuning strategy, for GRPO training. The test set consists of 100 samples, which are sufficient to cover a diverse range of scenarios. For quantitative evaluation, we compare the identity consistency reward score (abbreviated as ID-Consistency) produced by the reward model. Additionally, we conduct a user study in which participants are asked to select videos with higher ID-Consistency between baseline and Identity-GRPO. The winning rate is then calculated based on user preferences. We did not use existing face similarity metrics like ArcFace Deng et al. (2019) in the MH-IPV task, as they correlate with non-identity-related factors across frames and fail to capture overall video consistency with the reference image and prompt.
@@
Method ID-Consistency Winning Rate ($\%$)
VACE-1.3B 2.606 24
VACE-1.3B+Identity-GRPO 3.099 76
Phantom-1.3B 3.809 37
Phantom-1.3B+Identity-GRPO 4.056 63
@@
Main Results. Figure 1 illustrates the performance of Identity-GRPO on the test set throughout the training process. Both models exhibited a clear upward trend during the GRPO training process. As detailed in Table 3, Identity-GRPO improved ID-Consistency of VACE-1.3B and Phantom-1.3B by up to 18.9% and 6.5%, respectively. The advantage of Identity-GRPO was also evident in the user study, achieving winning rates of 76% and 63%. Qualitative comparison results are provided in Figure 2. In several cases, the baseline model produced outputs that clearly mismatched the reference image, whereas Identity-GRPO consistently maintained high identity alignment.

### v3

@@ 4.2 GRPO Training
Experiment Setup. We utilized the reference images from the preference dataset we constructed, along with the corresponding prompts generated by our prompt finetuning strategy, for GRPO training. The test set consists of 100 samples, which are sufficient to cover a diverse range of scenarios. For quantitative evaluation, we compare the identity consistency reward score (abbreviated as ID-Consistency) produced by the reward model. To evaluate visual quality and text relevance, we use two evaluation metrics from OpenS2V Yuan et al. (2025a): AestheticScore christophschuhmann (2024) and GmeScore Zhang et al. (2024b). Additionally, we conduct a user study in which participants are asked to select videos with higher ID-Consistency between baseline and Identity-GRPO. The winning rate is then calculated based on user preferences.
@@
Method ID-Consistency ↑ Aesthetics ↑ GmeScore ↑ Winning Rate ↑
[…]
@@
Main Results. Figure 1 illustrates the performance of Identity-GRPO on the test set throughout the training process. Both models exhibited a clear upward trend during the GRPO training process. As detailed in Table 3, Identity-GRPO improved ID-Consistency of VACE-1.3B and Phantom-1.3B by up to 18.9% and 6.5%, respectively. The advantage of Identity-GRPO was also evident in the user study, achieving winning rates of 76% and 63%. Furthermore, Identity-GRPO enhances the aesthetic scores of both VACE-1.3B and Phantom-1.3B, while the GME scores remain largely stable throughout the fine-tuning process. These results demonstrate that Identity-GRPO effectively strengthens identity preservation without compromising visual quality or prompt-following performance.
Qualitative comparison results are provided in Figure 2. In several cases, the baseline model produced outputs that clearly mismatched the reference image, whereas Identity-GRPO consistently maintained high identity alignment.
@@
Leveraging our curated high-quality dataset, we conducted a comparative analysis of two post-training methodologies: Supervised Fine-Tuning (SFT) and Identity-GRPO. To evaluate identity preservation, we employed the ArcFace similarity score as metrics. As shown in Table 5, the baseline VACE-1.3B achieved a score of 0.235, and SFT reached 0.261, whereas our proposed Identity-GRPO yielded a superior score of 0.298. These results demonstrate that Identity-GRPO significantly outperforms SFT in maintaining identity consistency, validating its efficacy in enhancing the generation quality of video models when trained on high-quality datasets.
@@
Table 5: Ablation study on different post-training methodologies. Bold: Best Performance.
VACE-1.3B Supervised Fine-Tuning Identity-GRPO
ArcFace similarity 0.235 0.261 0.298
ID-Consistency 2.606 2.774 3.099